OpenStudy (anonymous):

Find the derivative: d/dt (0to Sin t)∫[(1/(4-u^2)][du] NOTE: What I mean with o to Sin t is that the 0 is below the integral and sin t is up or above the integral.

OpenStudy (anonymous):

the derivative of the integral is the integrand, plus chain rule

OpenStudy (anonymous):

in simple english, take the integrand, replace \(u\) by \(\sin(x)\) and then multiply the result by the derivative of \(\sin(x)\) which is \(\cos(x)\)

OpenStudy (anonymous):

for a final answer I got cost/(4-(sint)^2)

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

Awesome! Thank you!

OpenStudy (anonymous):

yw (easy when you know what to do, right?)

OpenStudy (anonymous):

Sure is XP